find the surface area of a cylinder with a base diameter of 12 yd and a height of 7 yd. write your answer in terms of pie, and be sure to include the correct unit.
Answer by checkley77(12844)

SA=2*B*H
SA=2*2PIR/2*7
SA=2*2PI*12/2*7
SA=2*2PI*6*7
SA=2*12PI*7
SA=24*PI*7
SA=168*PI YDS^3 ANS.